### v1

# STG-Mamba: Spatial-Temporal Graph Learning via Selective State Space Model

Abstract
Spatial-Temporal Graph (STG) data is characterized as dynamic, heterogenous, and non-stationary, leading to the continuous challenge of spatial-temporal graph learning. In the past few years, various GNN-based methods have been proposed to solely focus on mimicking the relationships among node individuals of the STG network, ignoring the significance of modeling the intrinsic features that exist in STG system over time. In contrast, modern Selective State Space Models (SSSMs) present a new approach which treat STG Network as a system, and meticulously explore the STG system’s dynamic state evolution across temporal dimension. In this work, we introduce Spatial-Temporal Graph Mamba (STG-Mamba) as the first exploration of leveraging the powerful selective state space models for STG learning by treating STG Network as a system, and employing the Graph Selective State Space Block (GS3B) to precisely characterize the dynamic evolution of STG networks. STG-Mamba is formulated as an Encoder-Decoder architecture, which takes GS3B as the basic module, for efficient sequential data modeling. Furthermore, to strengthen GNN’s ability of modeling STG data under the setting of SSSMs, we propose Kalman Filtering Graph Neural Networks (KFGN) for adaptive graph structure upgrading. KFGN smoothly fits in the context of selective state space evolution, and at the same time keeps linear complexity. Extensive empirical studies are conducted on three benchmark STG forecasting datasets, demonstrating the performance superiority and computational efficiency of STG-Mamba. It not only surpasses existing state-of-the-art methods in terms of STG forecasting performance, but also effectively alleviate the computational bottleneck of large-scale graph networks in reducing the computational cost of FLOPs and test inference time.
1 Introduction
Spatial-temporal graph data is a kind of Non-Euclidean data which widely exists in our daily life, such as urban traffic network, metro system InFlow/OutFlow, social networks, regional energy load, weather observations, etc. Owning to the strong ability of capturing dynamic features across both temporal and spatial dimension, attention-based methods, such as Spatial-Temporal Transformer and its variants have recently been broadly implemented in spatial-temporal graph (STG) learning tasks. In the domain of STG traffic forecasting, seminal works have emerged that leverage these methods to enhance prediction accuracy. ASTGCN Guo et al. (2019) introduces an innovative approach combining spatial-attention & temporal attention mechanisms with graph neural networks. ADCT-Net Kong et al. (2024) presents to incorporate a dual-graphic cross-fusion Transformer for dynamic traffic forecasting. Huo et al. (2023) integrates hierarchical graph convolutional networks with Transformer networks to capture sophisticated STG relationships among traffic data. For weather forecasting, Liang et al. (2023) exemplifies how Transformer-based models can be utilized for accurate air quality forecasting across vast geographical areas. Chen et al. (2023) further demonstrates the power of GNNs in predicting air quality, emphasizing the importance of group-aware mechanisms in enhancing STG forecasting performance. In the realm of Social Recommendation, Huang et al. (2021) introduces a coupled GNN framework that incorporates inter-dependent knowledge graphs across both items and users. Wang et al. (2020a) explores the use of disentangled representations within a heterogeneous graph attention network to refine recommendation systems.
Although Transformer-based methods demonstrate notable improvements in STG relationship learning Chen et al. (2022); Cong et al. (2021), their modeling ability on large-scale STG networks and Long-Term Sequential Forecasting (LTSF) tasks is greatly hindered by the quadratic computational complexity $O(n^{2})$ provided by its core component–attention mechanism Zhou et al. (2021, 2022); Rampášek et al. (2022). Moreover, attention mechanisms usually approximate the full-length historical data, or encode all the context features, which may not be necessary for modeling the long-term dependencies of STG learning tasks. As such, the research world has been actively looking for counterparts of Attention-based models.
Lately, with the popularization of Mamba Gu and Dao (2023); Wang et al. (2024); Liu et al. (2024), Selective State Space Models (SSSM) have aroused considerable interest among researchers from computer vision and Natural Language Processing. SSSM is a variant of the State Space Models, which originated from Control Science and Engineering fields Lee et al. (1994); Friedland (2012) with strong theoretical foundation. State Space Models provide a professional framework to describe a physical system’s dynamics state evolution through sets of input, output, and state variables related by first-order differential or difference equations. The approach allows for a compact way to model and analyze systems with multiple inputs and outputs (MIMO) Aoki (2013).
As a pioneer deep learning-based SSSM, Mamba is proposed to address data-centric context learning for sequential tasks. Different from attention score calculation, Mamba construct selective state space models for the research objects, which encode contexts by using recurrent hidden states. The selection mechanism within Mamba allows it to control which part of the input data should be flowed into hidden states, as the significant context features that will influence the next-step embedding update.
STG learning can be treated as a complex process of understanding and forecasting the evolution of STG networks over time, which highly resembles the state space transition process where each state encapsulates the current configuration of a system and transitions represent changes over time. Deep learning-based SSSMs bring new horizons to STG learning tasks. However, great challenges pose in accurately and effectively adopting Mamba architectures for STG modeling, which is characterized as heterogeneous, dynamic, and non-stationary.
Motivated by the excellent long-term contextual modeling ability and low computational overhead of SSSMs, we propose Spatial-Temporal Graph Mamba (STG-Mamba) as the first deep learning-based SSSM for effective data-centric STG learning. The main contributions of this work are summarized in the following:
  • We make the first attempt to adapting SSSMs for STG learning tasks. A simple yet elegant way is employed to extend selective state space models to handle STG data. Specifically, we formulate the framework as Encoder-Decoder architecture, with $N$ stacked GS3B as basic modules and the proposed Kalman Filtering Graph Neural Network (KFGN) within it, to facilitate efficient Seq2Seq modeling.
  • A novel Graph Selective State Space Block (GS3B) is proposed for STG network pioneering integration with selective state space models, which performs input-dependent edge construction and dynamic node feature selection. GS3B is capable of capturing the sophisticated heterogenous dependencies among STG network, and accurately modeling the dynamic evolution of STG system over time, making it highly fitted for STG learning.
  • We introduce Kalman Filtering Graph Neural Network (KFGN) as the specialized adaptive spatiotemporal graph generation and upgrading method, which smoothly fits in within the SSSM-based context, ensuring the synchronous updates between graph structure and STG systems’s current state.
  • Extensive evaluations are carried on three open-sourced benchmark STG dataset. Results demonstrate that our STG-Mamba not only exceed other benchmark methods in terms of STG prediction performance, but also achieve $O(n)$ computational complexity, remarkably reducing the computational overhead compared with Transformers-based methods.
2 Preliminaries
Spatial-Temporal Graph (STG). STG is a kind of non-Euclidean structured data, which extensively existed in daily-life, business, and society. For a given graph network, if the node individuals have certain spatial connectivities at any given time step $\tau$, and the node features and/or graph structures are dynamically evolving over temporal axis, it could be defined as an STG network. Formally, we use $\mathbb{G}^{ST}=(V^{ST},E^{ST},A^{ST})$ to represent a given STG network. Here, $V^{ST}$ denotes the set of graph vertices (nodes) within the network, $E^{ST}$ denotes the edges of the graph network. $A^{ST}$ is the graph adjacency matrix, which indicates the calculated proximity between node individuals.
Spatial-Temporal Graph System (STG system). We for the first time define a network which consists of STG data as a Spatial-Temporal Graph System (STG system). In the theory foundation, a system is defined as the representation of a physical process, comprising state variables that describe the system’s current condition, input variables influencing the system’s state, and output variables reflecting the system’s response. For STG System, this framework is adapted to accommodate spatial dependencies and temporal evolution, structuring the system with nodes representing spatial entities, edges denoting spatial interactions, and state variables evolving over time to capture the dynamics of STG data.
State Space Models. State Space Models (SSMs) are a foundational concept in control science and system identification, offering a comprehensive framework to describe the dynamic evolutions of systems through mathematical theories. SSMs articulate a system’s behavior by defining a set of equations that represent the system’s state evolution over time, incorporating both control inputs and outputs.
Formally, SSMs are expressed through state equations $x_{t+1}=f(x_{t},u_{t})$ and observation equations $y_{t}=g(x_{t},u_{t})$, where $x_{t}$ denotes the state at time $t$, $u_{t}$ represents the control input, and $y_{t}$ is the observed output. The state transition function $f(\cdot)$ describes how the current state $x_{t}$ evolves to the next state $x_{t+1}$ under the influence of the control input $u_{t}$. $f(\cdot)$ captures the dynamics of the system, accounting for how internal and external factors modify the system’s state over time. The observation function $g(\cdot)$ is used for mapping the current state $x_{t}$ and control input $u_{t}$ to the observed output $y_{t}$. This function defines how the underlying state of the system is reflected in the measurements or observations collected.
As a variant of SSMs, Selective State Space Model emphasizes how to build the selection mechanism upon SSMs, which is highly resemble to the core idea of attention mechanism, making it a competitive counterpart of Transformer architecture.
STG Forecasting based on SSSMs. When employing Selective State Space Models (SSSMs) for spatial-temporal graph forecasting, the problem can be formulated as dynamically identifying and utilizing relevant portions of historical spatiotemporal data and graph structures to predict the future states of a STG system.
Given spatiotemporal graph $\mathbb{G}^{ST}=(V^{ST},E^{ST},A^{ST})$ and historical spatial-temporal sequential data $\rm{X}_{t-p+1:t}=\{X_{t-p+1},X_{t-p+2},...,X_{t}\}$, the objective is to utilize SSSMs to forecast future STG system’s states $\rm{\hat{X}_{t+1:t+k}}=\{\hat{X}_{t+1},...,\hat{X}_{t+k}\}$. This process is achieved by learning a mapping function $F_{SSSM}(\cdot)$ that dynamically selects relevant state transitions and interactions for prediction:
$F_{SSSM}(\rm{X}_{t-p+1:t};\mathbb{G}^{ST})=\rm{\hat{X}_{t+1:t+k}}$ (1)
3 Methodology
In this section, we start with introducing the model architecture of STG-Mamba and the formulation of STG Selective State Space Models. Following that, we elaborate the key contributions of this work, including the Kalman Filtering Graph Neural Networks (Section 3.2) which specifically designed for SSSMs to enable dynamic adaptive graph learning; and the Graph Selective State Space Blocks (GS3B) which acts as the basic modules for STG-Mamba (Section 3.3). Finally, we describe the computational efficiency of STG-Mamba in Section 3.4.
3.1 Formulation of STG Selective State Space Models
SSSMs like structured state space sequential models (S4) and Mamba, initially designed for sequential data, are adapted for Spatial-Temporal Graph (STG) data modeling by leveraging their structure to handle spatiotemporal dependencies dynamically. The adaptation involves discretizing the continuous system described by linear Ordinary Differential Equations (ODEs) into a format amenable to deep neural networks modeling and optimization, enhancing computational efficiency while capturing the evolving dynamics of the STG networks.
Figure 1: The comprehensive illustration of STG-Mamba’s Model Architecture.
Modern State Space Models all rely on a classical continuous system that maps a one-dimensional input function or sequence, denoted as $x(t)\in\mathbb{R}$, through intermediate implicit states $h(t)\in\mathbb{R}^{N}$ to an output $y(t)\in\mathbb{R}$. The aforementioned process can be represented as a linear Ordinary Differential Equation (ODE) as in Eq. 2:
$\begin{array}[]{lr}h^{\prime}(t)=\mathbb{A}h_{t}+\mathbb{B}x(t)\\ y(t)=\mathbb{C}h(t)\end{array}$ (2)
where $\mathbb{A}\in\mathbb{R}^{N\times N}$ represents the State Matrix, while $\mathbb{B}\in\mathbb{R}^{N\times 1},\mathbb{C}\in\mathbb{R}^{N\times 1}$ denote the Input Matrix and Output Matrix, respectively.
Obtaining the output sequence $y(t)$ at a given time-step $t$ necessitates determining $h(t)$ first, which poses great analytical challenges. As such, modern selective state space models address this challenge by converting the continuous system into a discretized version, enhancing the suitability for deep learning scenarios while boosting computational efficiency. A critical step in the discretization process involves the introduction of a time-scale factor $\Delta$, leading to the discrete representation of matrices $\mathbb{A}$ and $\mathbb{B}$ as $\overline{\mathbb{A}}$ and $\overline{\mathbb{B}}$, respectively, following a pre-determined discretization method, commonly the zero-order hold (ZOH) approach.
$\begin{array}[]{lr}\overline{\mathbb{A}}=\exp(\Delta\mathbb{A})\\ \overline{\mathbb{B}}=(\Delta\mathbb{A})^{-1}(\exp(\Delta A)-\mathbb{I})\cdot\Delta\mathbb{B}\end{array}$ (3)
After discretization, SSM-based methods can be computed in two ways: linear recurrent or global convolution, defined as in Eq. 4 and Eq. 5, respectively:
$\begin{array}[]{lr}h^{\prime}(t)=\overline{\mathbb{A}}h_{t}+\overline{\mathbb{B}}x(t)\\ y(t)=\mathbb{C}h(t)\end{array}$ (4)
$\begin{array}[]{lr}\overline{\mathbb{K}}=(\mathbb{C\overline{B}},\mathbb{C\overline{AB}},...,\mathbb{C\overline{A}^{L-1}\overline{B}})\\ y=x*\overline{\mathbb{K}}\end{array}$ (5)
where $\overline{K}\in\mathbb{R}^{L}$ denotes a structured convolutional kernel, and $L$ denotes the length of the input sequence $x$.
Figure 1 illustrates the comprehensive architecture of the proposed STG-Mamba. Specifically, we formulate the overall architecture following the Encoder-Decoder fashion for efficient sequential data modeling and prediction. In STG-Mamba, we leverage the Graph Selective State Space Block (GS3B) as the basic module, and repeat it $N$ times for Encoder and Decoder part. GS3B consists of several operations, including Layer Norm, Linear Layer, Kalman Filter Graph Neural Networks (KFGN), Graph State Space Layer (GSS-Layer), and Residual Connections. We specifically design KFGN and GSS-Layer for STG learning task build upon selective state space models.
3.2 Kalman Filtering Graph Neural Networks
Due to the dynamic, heterogenous, and non-stationary characteristics of STG data, the spatiotemporal dependencies of an STG system also evolve over time. When reflecting such graph dynamics in GNN formulation, it means the time-variant dynamic graph adjacency matrix. In existing literature, researchers usually employ various kinds of Attention-based methods to construct dynamic graph adjacency matrices, which may not be feasible in the context of state space model-based STG learning. On the other hand, Kalman Filtering (KF), as a well-known application of SSMs, has been extensively used in sequential data filtering and forecasting.
In light of that, we propose a novel approach which employs KF-based method to generate dynamic-evolving graphs, which delicately captures the system’s state-space transition for a given STG network. The method is termed as Kalman Filtering Graph Neural Networks (KFGN). To generate graph adjacency matrix $\mathbb{\alpha_{t}}$ at each time step using KFGN, the modeling procedure involves two steps as shown in the following:
Step 1: State Equation Calculation: Let $\mathbb{\alpha_{t}}$ denotes the graph adjacency matrix at time $t$, which indicates the current state of the STG system. The state equation is defined as $\mathbb{\alpha_{t}}=F_{t-1}\mathbb{\alpha_{t-1}}+W_{t-1}$, where $F_{t-1}$ is the state transition matrix that reflects changes in the graph adjacency matrix from $t-1$ to $t$, and $W_{t-1}$ represents the noise of the KF transition procedure.
Step 2: Observation Equation Calculation: The observed STG feature can be modeled as $X_{t}^{{}^{\prime}}=H_{t}\mathbb{\alpha_{t}}+V_{t}$, where $X_{t}^{{}^{\prime}}$ is the observation of STG feature at time $t$, $H_{t}$ is the observation matrix, and $V_{t}$ represents observation noise.
Iteratively applying these two equations through Kalman Filter enables dynamic estimation of the graph adjacency matrix $\alpha_{t}$ at each moment, adaptively capturing the evolving dynamics of the STG network. In practice, we employ a Neural Network-based method called Deep Kalman Filter (DKF) to automatically learn the state transition matrix $F_{t-1}$ and observation matrix $H_{t}$. Specifically, we adopt Gated Recurrent Unit (GRU) network as the architecture of Deep Kalman Filter. The GRU-based DKF is employed to learn the non-linear mapping for state transition and observation generation. The state transition matrix $F_{t-1}$ is modeled by GRU inner process to capture the non-linear relationships between states, predicting the state for the next moment. Similarly, the observation matrix $H_{t}$ is predicted under the current state by GRU. This process enables the system to automatically learn the dynamics of state transitions and observation generation from data without manually settings.
After receiving the $\alpha_{t}$, the next step of KFGN is the graph convolution operation, which is defined as:
$h_{t}^{kfgn}=(\omega_{gc}\odot\alpha_{t})x_{t}^{{}^{\prime}}$ (6)
where $\odot$ is the Hadamard product, and $x_{t}^{{}^{\prime}}$ denotes the input feature of KFGN at time $t$. $\omega_{gc}$ is the learnable weight matrix of KFGN.
3.3 GS3B: Graph Selective State Space Blocks
Inspired from Mamba, the GS3B is the major component of STG-Mamba, as depicted in Figure 1. After being processed by Layer Normalization, the input is divided into two sub-branches. In the first branch, the input STG embedding passed through a Linear Layer, followed by the proposed KFGN network and an activation function, before entering the GSS-Layer for selective feature learning. After that, the learned features are normalized by the Layer Norm. For the second branch, the input STG embedding undergoes a Linear Layer and an activation function. Subsequently, the two branches are fused together by an element-wise dot product operation. Finally, the merged embedding passes through a linear layer, and the outcome is combined with a residual connection to form the GS3B block’s output. In this research, Sigmoid Linear Unit (SiLU) is employed as the default activation function.
The GSS-Layer is a graph selective state space module specifically designed for STG data. GSS-Layer incorporates a selective mechanism which is capable of adjusting the SSM model’s parameters according to the input features. The characteristics enable the model to distinguish and retain key information which contributes significantly in STG learning tasks, while filtering out the irrelevant information. The pseudo-code of GSS-Layer (graph selective state space algorithm) is elaborated in Algorithm 1.
Algorithm 1 Pseudo Code of the Proposed Graph Selective State Space Algorithm
Input: $X\in\mathbb{R}^{batch\times L\times dim}$, the historical ground-truth STG data, where $batch$ is batch size of model training, $L$ is the historical time window length, $dim$ is the feature dimension of STG data;
$\mathbb{\alpha_{t}}$, the graph adjacency matrix of the current time step calculated by Deep Kalman Filter Network.
Params: $\mathbb{A}$ is the network param: nn.Parameter; $\mathbb{D}$ is the network param: nn.Parameter;
Operator: $Linear(\cdot)$, the linear projection layer;
Output:$y$, the processed feature with shape $\mathbb{R}^{batch\times L\times dim}$;
1:  $\Delta,\mathbb{B},\mathbb{C}=Linear(X),Linear(X),Linear(X)$
2:  $\overline{\mathbb{A}}=\exp(\Delta\mathbb{A})$
3:  $\overline{\mathbb{B}}=(\Delta\mathbb{A})^{-1}(\exp(\Delta\mathbb{A})-I)\cdot\Delta B$
4:  $h_{t}=(\overline{\mathbb{A}}\odot\mathbb{\alpha_{t}})h_{t-1}+\overline{\mathbb{B}}x_{t}$
5:  $y_{t}=\mathbb{C}h_{t}+\mathbb{D}x_{t}$
6:  $y=[y_{1},y_{2},...,y_{t},...,y_{L}]$.
7:  return $y$
3.4 Computational Efficiency of STG-Mamba
GS3B’s selection mechanism is introduced in Algorithm 1, which outlines the graph selective state space algorithm for efficiently processing STG data. It starts with initializing linear projections of the input STG data to obtain parameters $\Delta,\mathbb{B},\mathbb{C}$, which are used to compute the modified state transition matrix $\overline{\mathbb{A}}$ and input matrix $\overline{\mathbb{B}}$. These matrices incorporate the graph adjacency matrix $\mathbb{\alpha_{t}}$, updated by a Deep Kalman Filter Network, to calculate the hidden state $h_{t}$ for each time step. Finally, the output $y_{t}$ is generated by applying linear transformations to the hidden state and the input data, aggregating the results over the historical time window to form the output feature $y$. This algorithm effectively integrates dynamic graph structure information into SSSM, enhancing its capability to model sophisticated STG dynamics.
Leveraging a data-specific selection process, the increasing of sequence length $L$ or graph network node scale $N$ by $k$-fold can results in huge expansion of model parameters for Transformer-based methods. STG-Mamba introduces a sophisticated algorithm that is conscious of hardware capabilities, exploiting the layered structure of GPU memory to minimize the ensuing cost. In detail, for a batch size of $B$, STG feature dimension of $d^{\prime}$, and network node number of $N$, STG-Mamba reads the $O(BLd^{\prime}+Nd^{\prime})$ input data of $\mathbb{A},\mathbb{B},\mathbb{C},\Delta$ from High Bandwidth Memory (HBM), processes the intermediate stages sized $O(BLd^{\prime}N)$ in Static Random-Access Memory (SRAM), and dispatches the ultimate output, sized $O(BLd^{\prime})$, back to HBM. In contrast, if a Transformer-based model is employed, the HBM will read $O(BL^{2}d^{\prime})$ input data, and the computational size remains $O(BL^{2}d^{\prime})$ in SRAM. In common practices we often have $BL^{2}d^{\prime}\gg BLd^{\prime}+Nd^{\prime}$, which means SSSM-based approach can effectively decrease the computational overhead in large-scale parallel processing, such as neural network training. Additionally, by not retaining intermediate states, it also diminishes memory usage, since these states are recalculated during the gradient assessment in reverse phase. SSSM’s GPU-optimized design ensures STG-Mamba operates with a linear time complexity $O(L)$ relative to the length of the input sequence, marking a substantial speed advantage over traditional transformers’ dense attention computations, which exhibit quadratic time complexity $O(L^{2})$.
4 Experiments
4.1 Dataset Statistics and Baseline Methods
Datasets. For experimental evaluation, we meticulously select three real-world STG datasets from California road network speed records (PeMS04) Song et al. (2020), Hangzhou urban metro system entry/exit records (HZMetro) Liu et al. (2022), and weather station records across mainland China (KnowAir) Wang et al. (2020b). Acquired from the Caltrans Performance Measurement System (PeMS), PeMS04 records the average traffic speed from a total number of 307 sensors distributed on California Highway system. The time range of PeMS04 is 1st/Jan/2018 - 28th/Feb/2018, with 5-minute time interval. HZMetro was created with the In/Out transaction records of 80 stations of Hangzhou Metro system from 1st Jan 2019 to 25th Jan 2019, which takes 15-minute as data collection time interval. KnowAir documents weather observations in every 3-hour, encompassing a dataset that spans from 2015 to 2018. Within the KnowAir dataset, there are three subsets. In this study, we exclusively use subset No. 3, covering the period from 1/Sep/2016 to 31/Jan/2017. The detailed dataset statistics and descriptions are summarized in Table 1.
Table 1: Statistics and descriptions of the three datasets used for experimental evaluation.
Dataset PeMS04 HZMetro KnowAir
City California, USA Hangzhou, China 184 main cities across China
Data Type Network traffic speed Station InFlow/OutFlow Weather records
Total Nodes 307 80 184
Time Interval 5-minute 15-minute 3-hour
Time Range 1/Jan/2018 - 28/Feb/2018 1/Jan/2019 - 25/Jan/2019 1/Sep/2016 - 31/Jan/2017
Dataset Length 16,992 2,400 1,224
Baseline Methods. For fair comparison, we consider both benchmark spatiotemporal graph neural network (STGNN)-based methods, and Transformer-based methods for STG learning.
STGNN-based Methods:
(I) STGCN Yu et al. (2018): STGCN propose to capture the complex spatial-temporal dependencies within traffic network by graph neural networks, and design a fully convolutional approach for fast training.
(II) STSGCN Song et al. (2020): STSGCN presents a novel model which efficiently handling complex localized spatial-temporal correlations and data heterogeneity through a synchronous mechanism.
(III) STG-NCDE Choi et al. (2022): STG-NCDE leverages the concept of Neural Controlled Differential Equations (NCDEs) for spatial-temporal processing within a unified framework, achieving superior accuracy across six benchmark datasets.
(IV) DDGCRN Weng et al. (2023): DDGCRN dynamically captures the spatial-temporal features by separating the normal and abnormal signals and employing a novel graph convolutional recurrent network approach.
Attention (Transformer)-based Methods:
(V) ASTGCN Guo et al. (2019): ASTGCN leverages an attention-based spatial-temporal graph convolutional network to accurately predict STG data by dynamically capturing complex spatiotemporal correlations and patterns.
(VI) ASTGNN Guo et al. (2021): Attention based Spatial-Temporal Graph Neural Network incorporates a unique trend-aware self-attention and dynamic graph convolution to address temporal dynamics, spatial correlations, and the inherent periodicity and heterogeneity of STG data.
(VII) PDFormer Jiang et al. (2023): PDFormer introduces a tailored Transformer model, adept at capturing dynamic spatial dependencies and addressing the time delay in traffic condition propagation with its unique spatial self-attention and traffic delay-aware feature transformation.
(VIII) STAEformer Liu et al. (2023): STAEformer introduces a novel approach that enhances vanilla Transformer with a spatiotemporal adaptive embedding component, enabling it to capture complex STG patterns.
(IX) MultiSPANS Zou et al. (2024): MultiSPANS leverages multi-filter convolution modules, Transformers, and structural entropy optimization to effectively model complex multi-range dependencies in STG forecasting.
4.2 Implementation Settings
We split all the datasets with a ratio of 6:2:2 for Training/Validation/Testing sets along the time axis, respectively. Before starting model training session, all the data samples are normalized into range $[0,1]$ with MinMax normalization method. We set the batch size as 32 and employ AdamW as the Optimizer. The learning rate is initialized as $1e^{-3}$, with a weight decay of $5e^{-2}$. CosineAnnealingLR Loshchilov and Hutter (2016) is adopted as the learning rate scheduler, with a maximum of 50 iterations, and a minimum learning rate $1e^{-5}$. Model training epochs is set as 300. The number of Encoder/Decoder Layer is set as $N=8$ and $N^{\prime}=4$, respectively. We use the historical ground-truth data from the past 12 steps to forecast one future time step. Mean Squared Error (MSE) is employed as the Loss Function for optimization. And finally, we employ RMSE, MAE, MAPE as the evaluation metrics. All experiments are conducted under one NVIDIA A100 80GB GPU for neural computing acceleration.
4.3 Results Evaluation and Comparison
4.3.1 Experimental Results Comparison on Three STG Forecasting Tasks
A comprehensive evaluation between STG-Mamba and the proposed baseline methods is conducted, and we report the RMSE/MAE/MAPE performance results in Table 2. Our STG-Mamba consistently outperforms other benchmark methods on the three tasks, except for the MAPE criteria of PeMS04 (Flow). In PeMS04 (Flow) forecasting task, STAEformer shows slightly better MAPE results than STG-Mamba. It can also be easily observed that Transformer-based methods which integrate GNNs inside the architecture always gain superior performance than sole GNN-based methods. For instance, PDFormer, STAEformer, and MultiSPANS always show better performance than STGCN, STSGCN, and STG-NCDE. The reason is that although GNN is an effective approach of modeling STG tasks, Transformer-based methods have stronger ability to capture local and long-term dynamic dependencies. Within the GNN-based methods, STG-NCDE and DDGCRN stand out for their competitive performance. STG-NCDE models continuous-time dynamics within STG networks through neural controlled differential equations (NCDE), offering a more precise representation of spatiotemporal feature changes over time compared to other discrete GNN-based methods.
As a new contender poised to challenge Transformer architecture, Mamba introduces an innovative approach by replacing the foundational Attention mechanism with modern selective state space model (SSSM). SSSM diverges from Attention by utilizing a continuous-time model that captures the dynamic evolution of spatial-temporal dependencies more naturally and efficiently. Unlike Attention, which computes weights across all pairs of input and output positions and can be computationally intensive, Mamba’s selective state space model directly captures the temporal evolution of features, enabling it to scale more gracefully with the sequence length and complexity of spatial-temporal data. Furthermore, we facilitate STG-Mamba with Deep Kalman Filtering-based adaptive spatiotemporal graph evolution, and the elaborately designed Graph Selective State Space Block (GS3B), making it inherit the advantages of modern SSSM, and highly suitable for STG learning tasks.
Table 2: Performance Evaluation and Comparison with Benchmark Baselines on PeMS04, HZMetro, KnowAir Datasets.
Model PeMS04 (Flow) HZMetro KnowAir
RMSE MAE MAPE RMSE MAE MAPE RMSE MAE MAPE
STGCN 35.55 22.70 14.59 34.85 21.33 13.47 20.64 15.27 12.92
STSGCN 33.65 21.19 13.90 33.24 20.79 13.06 20.25 15.22 12.69
STG-NCDE 31.09 19.21 12.76 32.91 20.75 12.88 20.08 14.93 11.87
DDGCRN 30.51 18.45 12.19 31.69 19.52 12.45 19.63 12.45 11.84
ASTGCN 35.22 22.93 16.56 34.36 21.12 13.50 20.37 15.07 12.11
ASTGNN 31.16 19.26 12.65 32.75 20.63 12.47 19.96 12.38 11.23
PDFormer 29.97 18.32 12.10 30.18 19.13 11.92 18.65 12.21 11.06
STAEformer 30.18 18.22 11.98 29.94 18.85 12.03 18.49 11.93 10.75
MultiSPANS 30.46 19.07 13.29 30.31 18.97 11.85 18.57 12.24 10.88
STG-Mamba 29.53 18.09 12.12 29.26 18.38 11.67 18.05 11.73 10.66
4.3.2 Computational Efficiency Evaluation
STG-Mamba offers substantial improvement in computational efficiency in addition to performance gain. In order to quantitively evaluate computational efficiency, we select the widely recognized Floating Point Operations (FLOPs) and Inference Time as criteria. For both of the two criteria, a smaller value indicates better model performance. STG-Mamba is compare with other Transformer-based benchmark methods: ASTGNN, PDFormer, and STAEformer, on the same testing set.
Table 3 showcases the inference time comparison on PeMS04 and KnowAir testing set. It can be seen that ASTGNN is relatively redundancy, with the lowest inference time among the compared Transformer-based methods. STG-Mamba gains 18.09% and 20.54% performance increase in inference time than the 2nd best performance model (STAEformer).
We further compare the computational FLOPs increment with accordance to the scale complexity of STG network. Specifically, we use the number of node individuals to represent the scale of STG network. In the experiments, we select different number of nodes from PeMS04 and KnowAir. We select $[50,100,150,200,250,300]$ and $[30,60,90,120,150,180]$ node individuals for PeMS04 and KnowAir dataset, respectively. STG-Mamba is compared with the benchmark STAEformer in every settings. Figure 2 visualizes the FLOPs evaluation under different STG scale complexity. We can easily find that STG-Mamba enjoys a linear increment in FLOPs, while STAEformer shows quadratic increment in FLOPs. At the initial small-scale setting, the FLOPs for STG-Mamba/STAEformer are quite similar. However, with the increasing of network complexity and scale, the difference becomes more significant. At the 300 node setting of PeMS04, the FLOPs for STG-Mamba and STAEformer is 36.85G and 92.49G (2.50 times more), respectively. At the 180 node setting of KnowAir, the FLOPs for STG-Mamba and STAEformer is 57.08G and 118.13G (2.07 times more), respectively.
Benefiting from the linear computational increment brought by modern selective state space models, STG-Mamba and alike models demonstrate stronger capability and computational efficiency for modeling large-scale STG data, which holds promise for addressing the computational overhead in future large-scale STG learning.
Table 3: Test inference time comparison with benchmark Transformer-based approaches.
Model Inference Time on PeMS04 Inference Time on KnowAir
ASTGNN 4.85 s 5.31s
PDFormer 3.93 s 4.28s
STAEformer 3.76 s 4.09s
STG-Mamba 3.08 s 3.25s
performance $\uparrow$ +18.08% +20.54%
Figure 2: (Computational Efficiency Eval) FLOPs comparison between STG-Mamba and STAEformer on PeMS04/KnowAir dataset under different STG node number settings.
4.4 Ablation Study
Figure 3: Model component analysis of STG-Mamba.
To investigate the effectiveness of each model component within STG-Mamba, we further design three kinds of model variants, and evaluate their forecasting performance on PeMS04, HZMetro, and KnowAir dataset. The detailed description of these model variants are introduced in the following:
  1. 1.
    STG-Mamba: The full STG-Mamba model without any modification.
  2. 2.
    STG-Mamba w/o KFGN: We replace the proposed Kalman Filtering Graph Neural Networks (KFGN) within STG-Mamba with the standard Graph Convolution Network (GCN).
  3. 3.
    STG-Mamba w/o GSS-Layer: The GSS-Layer with proposed Graph Selective State Space Algorithm is replaced by the standard selective state space module as introduced in Gu and Dao (2023).
  4. 4.
    Transformer w KFGN: The GS3B block is reformulated as Transformer block (similar to ASTGNN Guo et al. (2021)), which consists of attention mechanism and the proposed KFGN.
As illustrated in Figure 3, STG-Mamba w/o KFGN shows slight worse performance than the full model, demonstrating the necessity and suitability of employing Kalman filtering-based dynamic GNN for SSSM-based methods. Different from the mainstream methods which employ various attention mechanisms to upgrade STG network’s adjacency matrix, our KFGN is more consistent with the SSSM context, and at the same time keep the linear algorithm complexity. STG-Mamba w/o GSS-Layer performs even worse than w/o KFGN. This is because in our design, GSS-Layer will take the generated STG adjacency matrix $\alpha_{t}$ at this timestamp as input for selective state space algorithm’s better upgradation. Replacing GSS-Layer with the standard selective state space layer leads to a less suitable version for STG learning. Finally, when we change from Mamba-based architecture to Transformer-based architecture, there is a substantial performance drop among all the three tasks. This discovery proves the effectiveness of modern SSSM over attention-based designs.
5 Conclusions
In this work, we first propose a novel spatial-temporal graph (STG) learning model named STG-Mamba that leverages modern selective state space models (SSSM) for accurate and efficient STG forecasting tasks. STG-Mamba is formulated as Encoder-Decoder structure, with the Graph Selective State Space Block (GS3B) as basic module. GS3B’s selection method facilitates input-dependent edge construction and node feature selection, successfully integrates STG network with SSSMs. Through the Kalman Filtering Graph Neural Networks (KFGN), the STG network achieves a smooth and dynamic evolution process of spatiotemporal graph adjacency matrix upgrading, which align with the whole graph selective state space modeling process. Compared with Attention-based methods, our STG-Mamba is able to achieve linear-time complexity and substantial reduction in FLOPs and test inference time. We conduct extensive empirical studies on three open-sourced STG forecasting datasets from urban traffic, metro In/Out flow, and weather records. Experimental results demonstrate the consistent superiority of STG-Mamba over other benchmark methods in both forecasting performance and computational efficiency. We believe that STG-Mamba presents a brand new promising approach to general STG learning fields, offering competitive model performance and input-dependent context-aware feature selection under an affordable computational cost.